OpenStudy (anonymous):

Double angle formulas: cos 4x - cos 2x = 0

OpenStudy (anonymous):

I dont know what to do with the cos 4x, however I understand what to do with the cos 2x

OpenStudy (anonymous):

cos 4x - cos 2x = 0 cos 4x - (2cos^2 x - 1) = 0

OpenStudy (anonymous):

The cos 4x is a mystery to me

OpenStudy (jhannybean):

\[\cos(2x) = 2\cos^2(x)-1 \implies \cos(4x) = 2\cos^2(2x)-1\]

OpenStudy (jhannybean):

Does that help?/

OpenStudy (anonymous):

Oh ok. and then do I use the same identity again on 2cos^2(2x) -1?

OpenStudy (jhannybean):

somehow you're going to have to turn it into a quadratic.

OpenStudy (jhannybean):

No you just have to change \(\cos(4x)\) to turn it into a quadratic.

OpenStudy (jhannybean):

So you will replace \(\color{red}{\cos(4x)}\) with \(\color{red}{2\cos^2(2x)-1}\)\[\color{red}{2\cos^2(2x)} -\cos(2x)\color{red}{-1} = 0\]

OpenStudy (anonymous):

I factor this now?

OpenStudy (jhannybean):

Then you can make this an easier function to solve for by: \(\color{blue}{a=\cos(2x)}\)\[2\color{blue}{a^2}-\color{blue}a-1=0\]

OpenStudy (anonymous):

So I dont factor. I just do that?

OpenStudy (anonymous):

Or do I factor then do that

OpenStudy (jhannybean):

No, you have to factor. I just helped making it easier to factor by substituting.

OpenStudy (jhannybean):

Factor: \(2a^2-a-1=0\)

OpenStudy (anonymous):

(a-2)(a+1)

OpenStudy (anonymous):

Or (2a-2)(a+1)?

OpenStudy (anonymous):

No (2a-1)(a+1)

OpenStudy (jhannybean):

\[(2a-1)(a+1) = 2a^2 +2a-a-1 = 2a^2 +a-1 \ne 2a^2-a-1\]

OpenStudy (anonymous):

!! (2a+1)(a-1)

OpenStudy (jhannybean):

\[(2a+1)(a-1) = 2a^2 -2a+a-1 = 2a^2 -a-1 ~\checkmark\]

OpenStudy (jhannybean):

Now that we have factored it, resubstitute the \(a\)

OpenStudy (anonymous):

(2cos(x)+1)(cos(x)-1)

OpenStudy (anonymous):

Then (2cos(x)+1) = 0 and (cos(x)-1) = 0

OpenStudy (jhannybean):

No, tha was not the correct substitution.

OpenStudy (jhannybean):

Remember, \(a = \cos(2x)\)

OpenStudy (anonymous):

(2cos(2x)+1) = 0 and (cos(2x)-1)

OpenStudy (jhannybean):

\[(2\cos(2x) +1) = 0\]\[\cos(2x)-1=0\]

OpenStudy (jhannybean):

Yes.

OpenStudy (anonymous):

So now I use the double angle on both equations right?

OpenStudy (anonymous):

(4cos^2(x)-1+1) = 0 and (2cos^2(x)-1-1)

OpenStudy (anonymous):

So 4cos^2(x) = 0 2cos^2(x) -2 =0

OpenStudy (anonymous):

4cos^2(x) = 0 cos^2(x) = 1

OpenStudy (jhannybean):

Hmm, trying to figure it out.

OpenStudy (anonymous):

arccos(0) = pi/2 and 3pi/2 arccos(sqrt(1)) = 0

OpenStudy (anonymous):

?

OpenStudy (jhannybean):

I'll check my answer with yours in a sec,

OpenStudy (jhannybean):

\[2\cos(2x) +1 = 0\]\[2\cos(2x)=-1\]\[\cos(2x)=-\frac{1}{2}\]Take the inverse to solve for \(2x\)\[2x= \cos^{-1}\left(-\frac{1}{2}\right)\]

OpenStudy (jhannybean):

What radians give -1/2?

OpenStudy (anonymous):

7pi/6 and 11pi/6

OpenStudy (anonymous):

And then I multiply that by 2?

OpenStudy (jhannybean):

Not exactly.

OpenStudy (jhannybean):

cosine is negative in Q2.

OpenStudy (anonymous):

If I'm trying to find a negative x isn't cos negative?

OpenStudy (jhannybean):

and 3, but \(\frac{7\pi}{6} = -\dfrac{\sqrt{3}}{2}\)

OpenStudy (anonymous):

OH I did sine didnt I.

OpenStudy (jhannybean):

Yes.

OpenStudy (anonymous):

4pi/3 and 5pi/6

OpenStudy (jhannybean):

Yes and no.

OpenStudy (anonymous):

Cos cannot be negative?